(Ratio Analysis): Last year Co. XYZ had sales of $ 400,000, with “cost of goods sold” of $ 112,000. The firm had operating expenses of $ 130,000, and an increase in retained earnings of $ 58,000. There are currently 22,000 common shares issued and the firm pays $ 1.60 dividend per share.

A. Assuming that the firm's profits are taxed at 34%, build an “Income Statement” for the firm.

B. Compute the "Operating Profit Margin" for the signature.

C. What was the “times interest earned?
Answer #1

Dividends Paid = Dividend per share * No. of common shares outstanding

= $1.60 * 22,000 = $35,200

Net Income = Dividends Paid + Increase in retained earnings

= $35,200 + $58,000 = $93,200

EBT = Net Income / (1 - t) = $93,200 / (1 - 0.34) = $93,200 / 0.66 = $141,212.12

Income Statement

Sales $400,000

Less: Cost of goods sold ($112,000)

Less: Operating Costs ($130,000)

EBIT $158,000

Less: Interest ($ 16,787.88)

Earnings before tax   $141,212.12

Less: Taxes(@34%) ($ 48,012.12)

Net Income $ 93,200

b). Operating Profit Margin = EBIT / Sales = $158,000 / $400,000 = 0.395, or 39.50%

c). Times Interest Earned = EBIT / Interest Expense = $158,000 / $16,787.88 = 9.41